Sin+cos=1,find sin-cos ?
Answer by richard1234(7193)

Sine and cosine need an argument; the expressions sin = 1/2 or sin+cos = 1 make no sense.

Assuming you mean sin x + cos x = 1, sin x - cos x = ?, we have (sin x + cos x)^2 = 1 + 2 sin x cos x = 1 (using Pythagorean identity) so sin x cos x = 0.

Also (sin x - cos x)^2 = 1 - 2 sin x cos x = 1 - 2*0 = 1. So sin x - cos x could equal 1 or -1.